Fractions (Answers should be in simplest form)

1 ¹/² + ¾ =


1
Expert's answer
2022-03-08T05:15:02-0500

"1\frac{1} {2} +\frac{3}{4}=1+\frac{1 } {2 } +\frac{3}{4}="


"=\frac{4} {4} +\frac{2} {4 } +\frac{3}{4}=\frac{9} {4} =\frac{8} {4} +\frac{1} {4 }="


"=2+\frac{1 }{4} =2\frac{1} {4}"